# Can World Models Achieve Mapless Path Planning with Semantic Targets?

### v1

@@ Abstract
While recent world models generate highly realistic videos, their ability to perform robot path planning remains unclear and unquantified. We introduce Target-Bench, the first benchmark specifically designed to evaluate world models on mapless path planning toward semantic targets in real-world environments. Target-Bench provides 450 robot-collected video sequences spanning 45 semantic categories with SLAM-based ground truth trajectories. Our evaluation pipeline recovers camera motion from generated videos and measures planning performance using five complementary metrics that quantify target-reaching capability, trajectory accuracy, and directional consistency. We evaluate state-of-the-art models including Sora 2, Veo 3.1, and the Wan series. The best off-the-shelf model (Wan2.2-Flash) achieves only 0.299 overall score, revealing significant limitations in current world models for robotic planning tasks. We show that fine-tuning an open-source 5B-parameter model on only 325 scenarios from our dataset achieves 0.345 overall score—an improvement of more than 400% over its base version and 15% higher than the best off-the-shelf model.
Figure 1: Target-Bench provides a dataset collected with a quadruped robot, and a benchmark for evaluating world models in mapless path planning toward text-specified goals with implicit semantic meaning. In Target-Bench, world models receive a camera frame and a textual prompt describing the target state, and predict a future video depicting the trajectory toward the goal. A world decoder then extracts the planned path from this video, which is compared against the maneuver executed by a human-operated quadruped.
footnotetext: {*} Equal contribution; author order settled via Mario Kart.
@@ 1 Introduction
“If one day, we bring a humanoid robot to this conference venue, a place it has never seen before, and I can simply say ‘Please bring this bottle of water to someone in the audience,’ and it can smoothly walk over and do it by itself, I think that would be the robot’s ChatGPT moment.”
— Xingxing Wang, Head of Unitree, at the World Robot Conference 2025.
Embodied AI has been advancing rapidly, while its core challenges are becoming clearer. Despite major improvements (better actuators and sensors) in robotic hardware, AI software remains the main bottleneck limiting robots from realizing their full potential [long2025survey]. As a result, the robotics and embodied AI community is seeking to develop AI systems that are more robust and more generalizable. Recent breakthroughs in World Models (WMs) [ball2025genie3, veo3deepmind, wiedemer2025video, sora2, wan2025wan] have drawn attention from the community. WMs learn to predict how the world evolves over time [ha2018worldmodels, lecun2022path]. Given an initial observation (e.g., an image frame) and a condition (e.g., a text prompt or an action), models such as Veo3.1 [veo3deepmind] and Genie3 [ball2025genie3] can generate future frames with high spatio-temporal consistency. More importantly, they exhibit remarkable reasoning abilities in visual semantics and causal relationships [videoworldsimulators2024, wiedemer2025video, gao2025foundation]. This potential has sparked interest within the embodied AI community in applying WMs to robotic applications [li2025robotic, zhu2025unified, pang2025learning, wang2025enhancing]. Recent approaches from Unitree, such as UnifoLM-WMA-0 [unifolm], aim to bridge WMs with low-level robot control.
@@
The underlying philosophy of these approaches is that if a model can accurately predict how the world evolves, its predictions can serve as a plan to guide the robot actions. However, a key question remains: How accurate must these predictions be to count as useful for planning? More broadly, how can we quantitatively assess a WM ’s reasoning, task-solving, and planning ability? Existing evaluation frameworks focus mainly on reconstruction quality, from visual fidelity to physical consistency [li2025worldmodelbench, duan2025worldscore], while assessing mapless path planning driven by semantic understanding remains an open challenge [long2025survey]. Bridging this gap requires a benchmark that evaluates not just how realistic a WM’s predictions look, but whether they contain the semantic reasoning needed for robotic planning tasks.
@@
To address this challenge, we introduce Target-Bench. To our knowledge, it is the first benchmark evaluating WMs for mapless path planning toward semantic targets in unstructured real-world environments. The target state is specified by user text and can have implicit semantic meaning (Fig. 1). Our contributions are:
@@
  • An open-source dataset of 450 videos (112,500 frames) collected with a quadruped robot, covering 45 semantic target categories in diverse indoor and outdoor environments, with ground truth trajectories and human annotations for explicitly- and implicitly-stated targets.
  • The first systematic evaluation pipeline for WMs in mapless path planning with textual semantic goals.
  • A world decoder that extracts camera motion from generated videos, comparing spatial reconstruction methods and introducing a new scale recovery technique.
@@
  • A comparative study of open-source and proprietary WMs, including the first fine-tuning of an open-source model on a small real-world dataset for path planning, demonstrating improved generalization to unseen environments over proprietary models.
@@ 2 Related Work
Early approaches such as World Models [ha2018worldmodels] and Dreamer [hafner2020dreamcontrol] introduced latent state-space models, demonstrating that generative prediction has the potential to support planning and control [gao2025foundation]. More recently, foundation-scale methods pretrained on large video corpora have emerged: Cosmos [nvidia2025cosmos] compared diffusion and autoregressive paradigms; DIAMOND [alonso2024diamond] improved visual fidelity; V-JEPA 2 [assran2025vjepa2] enhanced efficiency by forecasting in latent space. Advances in video generation, including Sora 2 [sora2], Veo 3.1 [veo3deepmind, wiedemer2025video], and Wan [wan2025wan], have further improved prompt adherence and physics-aware dynamics, enabling long-horizon predictions. Interactive models such as Genie-1 [bruce2024genie1], Genie-2 [parkerholder2024genie2], and Genie-3 [ball2025genie3] provide controllable environments for use in robotics and games. Unitree’s recent UnifoLM-WMA-0 [unifolm] aims to extract the reasoning capabilities of WMs and translate them into real-world actions. However, the ability of WMs to perform robot task planning in real-world scenarios is largely untested.
@@
Recent works have proposed benchmarks to evaluate generative WMs, often relying on Vision Language Model (VLM)-based judges to assess the quality and consistency of generated scenes. VBench [huang2024vbench] targets 16 fine-grained video quality dimensions, while WorldModelBench [li2025worldmodelbench] focuses on physics consistency (e.g., Newtonian motion). Compared to VBench, WorldScore [duan2025worldscore] emphasizes controllability and dynamics in next-scene prediction. World-in-World [zhang2025world] enables closed-loop evaluation but uses WMs as simulators and does not test their planning abilities. Overall, existing benchmarks primarily assess spatio-temporal or physical consistency, and do not consider high-level robot planning. Recent advances in spatio-temporal scene reconstruction [wang2025vggtvisualgeometrygrounded, xiao2025spatialtrackerv2, huang2025vipe] further highlight the limits of using only VLM-based scoring for WM evaluation. We extend this literature by assessing not only spatio-temporal coherence but also path planning abilities.
(a) Quadruped robot hardware setup and communication network.
(b) SLAM pipeline for state estimation.
@@
(a) All trajectories.
(b) Word cloud of captions.
(c) Semantic target categories.
(d) Target-Bench Evaluation Results.
Figure 3: Visualization of dataset structure and semantics.
Semantic Navigation Datasets.
Several datasets address semantic robot navigation, with social compliance in real-world scenarios. EgoWalk [akhtyamov2025egowalk] provides egocentric data with stereo images, odometry, and traversability labels, but suffers from heading misalignment and IMU drift. LeLaN [hirose2024lelan] leverages unlabeled human walking videos to train language-conditioned policies, while SCAND [karnan2022socially] offers socially compliant robot demonstrations. MuSoHu [nguyen2023toward] captures multimodal wearable data for natural interactions, SACSoN [hirose2023sacson] focuses on office navigation with human–robot interaction, and SANPO [waghmare2025sanpo] targets outdoor navigation. CityWalker [liu2025citywalker] enables mapless urban navigation amid traffic and crowds. However, these datasets lack explicit navigation targets, and their VLM-based annotations are not target-oriented.
@@ 3 Target-Bench
As shown in Fig. 1, Target-Bench consists of two components: a Target dataset and the Target benchmark evaluation pipeline. The Target dataset (Fig. 2(a)) is collected with a quadruped robot equipped with multi-modal sensors in diverse indoor and outdoor environments. It spans 45 semantic target categories (e.g., doors, chairs, trees) and includes 450 scenarios. Our Target benchmark comprises two stages: world decoder and path evaluation. The world decoder extracts camera poses from generated video to form a path, which is then compared against the ground truth path. Path evaluation metrics focus on the proximity to the target and directional consistency, jointly assessing scenario reconstruction quality and semantic navigation capability.
@@ 3.1.1 Quadruped Robot Platform
As shown in Fig. 2(a), our data-collection platform is built on a DEEP Robotics Lite 3 Venture quadruped. It carries a Livox Mid-360 LiDAR, an OAK-D Pro W stereo RGB camera, and an NVIDIA Jetson AGX Orin for mapping and state estimation. The LiDAR connects to the robot base via an Ethernet switch, and the camera connects to the Jetson via USB. The Jetson is powered by a dedicated onboard battery. A Wi-Fi link connects the Jetson to a remote laptop for monitoring and logging, and to a handheld controller for teleoperation. Our software stack (Fig. 2(b)) uses a LiDAR-centric SLAM pipeline [segalGeneralizedICP2009, kümmerleg2o2011] with multi-sensor fusion. Inertial Measurement Unit (IMU) data and legged odometry are fused via an Extended Kalman Filter (EKF) to produce a stable base-frame pose, which is broadcast through the ROS TF tree to align the LiDAR and robot frames. The LiDAR point clouds are processed by the Simultaneous Localization and Mapping (SLAM) front-end with motion compensation and incremental registration guided by the fused odometry. The back-end then optimizes the trajectory and builds a global map, enabling accurate pose estimation.
@@ 3.1.2 Semantic Target Data Collection
Each data sample in the Target dataset consists of four components: a video sequence, a ground truth trajectory, a semantic target and a point cloud map. Video frames are captured at 25Hz, yielding approximately 10 seconds of continuous observation per sample. The semantic targets are annotated and pre-selected by human experts to ensure diversity and relevance for navigation tasks. Fig. 3(a) shows that our trajectories span diverse directions and movement patterns, providing balanced and realistic navigation coverage. Fig. 3(c) presents the semantic target distribution with a variety of scenarios. This mix captures both indoor and outdoor environments with common static and dynamic objects. Fig. 3(b) highlights the dataset’s rich caption vocabulary. The targets are defined as either explicit or implicit. Explicit targets specify the object name directly, whereas implicit targets describe the object through its attributes or functions embedded in the prompt context (Fig. 4). Overall, the whole dataset is divided into a train split (325 scenarios) and a benchmark (evaluation) split (125 scenarios).
@@ 3.2 Target Benchmark Architecture
The Target benchmark architecture (Fig. 5) consists of two main components designed to quantify world model planning capabilities. First, the world decoder extracts camera trajectories from generated videos using state-of-the-art 3D reconstruction methods. Second, the path evaluation module compares predicted trajectories against SLAM-verified ground truth using a comprehensive suite of metrics.
@@ 3.2.1 World-Decoder
As shown in Fig. 5, to extract camera trajectories from WM’s generated videos, we employ three state-of-the-art 3D reconstruction methods: VGGT [wang2025vggtvisualgeometrygrounded], SpaTracker [xiao2025spatialtrackerv2], and ViPE [huang2025vipe]. Each method processes video frames to recover camera poses, but differs in their approach and output characteristics.
VGGT.
Visual Geometry Grounded Transformer (VGGT) is a feed-forward transformer that directly predicts camera poses, depth maps, and point clouds from multi-view images. Given a sequence of $S$ images $\mathcal{I}=\{I_{1},I_{2},\ldots,I_{S}\}$, VGGT first encodes them through a vision transformer to obtain aggregated tokens. A camera head then predicts a pose encoding $\mathbf{p}_{s}=[\mathbf{T}_{s};\mathbf{q}_{s};\mathrm{\mathbf{fov}}_{s}]\in\mathbb{R}^{9}$ for each frame $I_{s}$, where $\mathbf{T}_{s}\in\mathbb{R}^{3}$ is the camera translation matrix, $\mathbf{q}_{s}\in\mathbb{R}^{4}$ is the rotation quaternion, and $\mathrm{\mathbf{fov}}_{s}\in\mathbb{R}^{2}$ represents the horizontal and vertical field of view. This encoding is converted to standard camera parameters $\mathbf{E}_{s}=[\mathbf{R}(\mathbf{q}_{s})\mid\mathbf{T}_{s}]\in\mathbb{R}^{3\times 4}$, where $\mathbf{R}(\mathbf{q}_{s})$ converts the quaternion to a rotation matrix.
SpaTracker.
SpaTracker extends VGGT with tracking for better temporal consistency. It adopts a two-stage design: (1) VGGT4Track predicts initial poses $\{\mathbf{E}_{s}^{(0)}\}_{s=1}^{S}$ and depths $\{D_{s}\}_{s=1}^{S}$; (2) a tracking module refines these poses via point correspondences across frames. Given query points $\mathbf{Q}\in\mathbb{R}^{N\times 3}$ on the first frame, the tracker predicts their 3D trajectories and optimizes camera poses through bundle adjustment:
$\{\mathbf{E}_{s}^{*}\}=\arg\min_{\{\mathbf{E}_{s}\}}\sum_{s=1}^{S}\sum_{n=1}^{N}\rho\left(\|\pi(\mathbf{K}_{s}\mathbf{E}_{s}\mathbf{X}_{n})-\mathbf{x}_{n,s}\|^{2}\right)$ (1)
where $\mathbf{X}_{n}$ is the 3D position of point $n$, $\pi(\cdot)$ the projection, $\mathbf{x}_{n,s}$ the tracked 2D point, and $\rho(\cdot)$ a robust loss. The result is a refined camera-to-world trajectory $\{\mathbf{C}_{2W,s}\}_{s=1}^{S}$, where each $\mathbf{C}_{2W,s}\in\mathbb{R}^{4\times 4}$ is the full camera-to-world transform. As in VGGT, SpaTracker outputs are at arbitrary scale.
ViPE.
ViPE is a SLAM-based visual-inertial pipeline providing metric-scale poses by fusing visual and inertial data. Given RGB frames and IMU inputs, it performs incremental pose estimation with loop closure. Unlike VGGT and SpaTracker, ViPE directly outputs paths in metric units, removing the need for scale recovery. The final path is a sequence of SE(3) transformations $\{\mathbf{T}_{s}\}_{s=1}^{S}$, where each $\mathbf{T}_{s}\in\mathbb{R}^{4\times 4}$ encodes rotation and translation in meters.
@@
Monocular methods such as VGGT and SpaTracker estimate camera motion only up to an unknown global scale. We restore metric consistency at the segment level by anchoring predictions to a single scalar scale factor $\lambda$ derived from ground truth displacement. Let $\mathbf{E}_{1},\mathbf{E}_{k}\in\mathbb{R}^{3\times 4}$ be the predicted extrinsic matrices for the first and the $k$-th frame. We extract the translation vectors $\mathbf{t}_{1}$ and $\mathbf{t}_{k}$ as the fourth column of $\mathbf{E}_{1}$ and $\mathbf{E}_{k}$. The predicted displacement is $d_{\mathrm{pred}}=\|\mathbf{t}_{k}-\mathbf{t}_{1}\|_{2}$, and the scale factor is $\lambda=d_{\mathrm{real}}/d_{\mathrm{pred}}$, while the real displacement is $d_{\mathrm{real}}$. We then rescale all predicted translations uniformly, $\mathbf{t}_{s}^{\mathrm{scaled}}=\lambda\,\mathbf{t}_{s}$ for $s=1,\ldots,S$. This lightweight, data-anchored normalization preserves relative geometry while lifting trajectories to meters, enabling direct and fair comparison against ground truth paths. Note that ViPE inherently produces metric-scale outputs via sensor fusion and therefore does not require this step.

### v2

# https://target-bench.github.io/

@@ Abstract
While recent video world models can generate highly realistic videos, their ability to perform semantic reasoning and planning remains unclear and unquantified. We introduce Target-Bench, the first benchmark that enables comprehensive evaluation of video world models’ semantic reasoning, spatial estimation, and planning capabilities. Target-Bench provides 450 robot-collected scenarios spanning 47 semantic categories, with SLAM-based trajectories serving as motion tendency references. Our benchmark reconstructs motion from generated videos with a metric scale recovery mechanism, enabling the evaluation of planning performance with five complementary metrics that focus on target-approaching capability and directional consistency. Our evaluation result shows that the best off-the-shelf model achieves only a 0.341 overall score, revealing a significant gap between realistic visual generation and semantic reasoning in current video world models. Furthermore, we demonstrate that fine-tuning process on a relatively small real-world robot dataset can significantly improve task-level planning performance.
Keywords:
World models Semantic reasoning Tasking benchmark
Figure 1: A Video World Model’s planning ability is evaluated by our benchmark. We provide a diverse dataset containing annotated semantic goal, task description and motion reference to establish this novel benchmark. The generated future visual predictions are processed by our world-decoder, and the results are then evaluated with meticulously designed metrics that focus on tendency rather than trajectory overlap.
*equal contribution
@@ 1 Introduction
Embodied AI has been advancing rapidly, while its core challenges are becoming increasingly clear. Despite major improvements in robotic hardware (e.g. actuators), the embodied intelligence remains the main bottleneck limiting robots from realizing their full potential [long2025survey]. As a result, the robotics community is seeking to develop embodied AI systems that are more robust and generalizable.
World Models (WMs) learn to predict how the world evolves over time [ha2018worldmodels, lecun2022path]. In early 2024, Sora [videoworldsimulators2024], a video generation model, was the first to self-claim as a world evolution simulation model. Recent breakthroughs in video models demonstrate remarkable reasoning abilities in visual semantics and causality [videoworldsimulators2024, wiedemer2025video, gao2025foundation]. In this regard, these video models can also be regarded as world models [ball2025genie3]. Given an initial observation (e.g., an image frame) and a condition (e.g., a text prompt), video WMs such as Veo 3.1 [veo3deepmind] and Sora 2 [sora2] can generate future frames with photorealistic quality and high spatio-temporal consistency.
@@
Unsurprisingly, video WMs have sparked growing interest from the robotics community [li2025robotic, zhu2025unified, pang2025learning, wang2025enhancing]. Recent approaches such as UnifoLM-WMA-0 [unifolm] aim to integrate video WMs with low-level robot control. The underlying philosophy of these approaches is that if a model can accurately predict how the world evolves, its predictions can serve as a plan to guide the robot actions. However, a key question remains: How accurate must these predictions be to count as useful for planning? More broadly, how can we quantitatively assess a video WM ’s reasoning, task-solving, and planning ability? Existing evaluation frameworks focus mainly on visual fidelity and spatio-temporal consistency [li2025worldmodelbench, duan2025worldscore], while assessing mapless path planning that require strong semantic understanding remains an elusive task [long2025survey]. Bridging this gap requires a benchmark that can holistically evaluate video world models’ semantic reasoning beyond visual quality and spatio-temporal consistency.
@@
To address this challenge, we introduce Target-Bench. To the best of our knowledge, it is the first benchmark for evaluating video WMs for mapless path planning toward semantic targets in unstructured real-world environments. The semantic targets are described in natural language, and many cases are designed to include implicit semantic meanings (Fig. 1). Our contributions are as follows:
  • The first systematic benchmark for evaluating video world models on mapless path planning with textual semantic goals, featuring a modular and replaceable world-decoder and a meticulously designed set of metrics, which makes comprehensive evaluation of video world models’ semantic reasoning, spatial estimation, and planning capabilities possible.
@@
  • An open-source dataset of 450 semantic-target-oriented scenarios (112,500 frames) collected with a quadruped robot, covering 47 categories across diverse indoor and outdoor environments, with ground truth trajectories and human annotations for explicitly and implicitly described targets.
@@
  • A comparative study of open-source and proprietary WMs, including the first fine-tuning of an open-source model on a small real-world robot dataset for path planning, showing that data-efficient adaptation can substantially boost task-level planning performance. We open-source all code and dataset.
@@ 2 Related Work
Early approaches such as World Models [ha2018worldmodels] and Dreamer [hafner2020dreamcontrol] introduced latent state-space models, demonstrating that generative prediction has the potential to support planning and control [gao2025foundation]. More recently, foundation-scale methods pretrained on large video corpora have emerged: Cosmos [nvidia2025cosmos] compared diffusion and autoregressive paradigms; DIAMOND [alonso2024diamond] improved visual fidelity; V-JEPA 2 [assran2025vjepa2] enhanced efficiency by forecasting in latent space. Advances in video WMs, including Sora 2 [sora2], Veo 3.1 [veo3deepmind, wiedemer2025video], and Wan [wan2025wan], have further improved prompt adherence and physics-aware dynamics, enabling long-horizon predictions. Interactive WMs such as Genie-1 [bruce2024genie1], Genie-2 [parkerholder2024genie2], and Genie-3 [ball2025genie3] provide controllable environments for use in robotics and games. Unitree’s recent UnifoLM-WMA-0 [unifolm] aims to extract the reasoning capabilities of WMs and translate them into real-world actions. However, the ability of WMs to perform robot task planning in real-world scenarios is largely untested.
Semantic Navigation Datasets.
Navigation datasets such as R2R [anderson2018vision] and RxR [ku2020room] are primarily restricted to indoor simulation environments. To address real-world complexity, datasets such as MuSoHu [nguyen2023toward], EgoWalk [akhtyamov2025egowalk], and SANPO [waghmare2025sanpo] capture multimodal data via wearable sensors, though the latter faces heading misalignment and IMU drift. Others utilize human videos or robot demonstrations for language-conditioned policies (LeLaN [hirose2024lelan]) and socially compliant interaction (SCAND [karnan2022socially], SACSoN [hirose2023sacson]). For broader coverage, Sekai [li2025sekai] and CityWalker [liu2025citywalker] leverage web-scale videos to achieve global diversity, although trajectories are only annotated up-to-scale. Crucially, as compared in Table 1, these datasets lack explicit navigation targets, and their annotations are descriptive rather than target-oriented.
@@
Evaluation of WMs has transitioned from assessing visual quality and camera controllability (VBench [huang2024vbench], WorldScore [duan2025worldscore]) and basic physics (WorldModelBench [li2025worldmodelbench]) to functional utility. While frameworks like World-in-World [zhang2025world] and WorldArena [shang2026worldarena] enable closed-loop testing, they are predominantly confined to simulation. Crucially, many recent robotics-oriented suites (RBench [deng2026rethinking], EWMBench [yue2025ewmbench], WorldEval [li2025worldeval]) rely on subjective Vision Language Model (VLM)-based judges for success detection, which introduces inherent judge uncertainty and limits result reproducibility. Moreover, existing benchmarks typically focus on explicit semantic instruction rather than implicit semantic reasoning. A concise comparison across these dimensions is provided in Table 2.
(a) Quadruped robot hardware setup.
(b) SLAM pipeline.
@@
(a) Trajectories.
(b) Caption word cloud.
(c) Data scenario environments.
(d) Semantic target categories.
Figure 3: Visualization of benchmark dataset structure and semantics: (a) spatial distribution of trajectories, (b) word cloud representation of caption vocabulary, (c) data collection environments, and (d) distribution of semantic target categories.
@@ 3 Target-Bench
As shown in Fig. 1, Target-Bench consists of two components: a Target dataset and the Target benchmark evaluation pipeline. The Target dataset (Fig. 2(a)) is collected with a quadruped robot equipped with multi-modal sensors in diverse indoor and outdoor environments. It spans 47 semantic target categories (e.g., cars, chairs, bicycles) and includes 450 scenarios. Our Target benchmark comprises two stages: world-decoder and path evaluation. The world-decoder extracts camera poses from generated video to form a path, which is then compared against the ground truth path. Path evaluation metrics focus on the proximity to the target and directional consistency, jointly assessing scenario reconstruction quality and semantic navigation capability.
Table 1: Comparison with existing navigation datasets. Our dataset is collected via target-oriented teleoperation of a real quadruped robot in real-world environments, providing video sequences paired with semantic target prompts, point cloud maps, and metric-scale trajectories with kinematic consistency.
Dataset Source Type Metric Traj. Semantic Target Point Cloud Kinem. Match
R2R [anderson2018vision] Simulation
RxR [ku2020room] Simulation
[…]
MuSoHu [nguyen2023toward] Real-Human
SACSoN [hirose2023sacson] Real-Robot
SANPO [waghmare2025sanpo] Sim./Real
Sekai [li2025sekai] Sim./Real
[…]
SCAND [karnan2022socially] Real-Robot
EgoWalk [akhtyamov2025egowalk] Real-Human
Ours Real-Robot
@@ 3.1.1 Quadruped Robot Platform
As shown in Fig. 2(a), our data-collection platform is built on a DEEP Robotics Lite 3 Venture quadruped robot. It carries a Livox Mid-360 LiDAR, an OAK-D Pro W stereo RGB camera, and an NVIDIA Jetson AGX Orin for mapping and state estimation. The LiDAR connects to the robot base via an Ethernet switch, and the camera connects to the Jetson via USB. The Jetson is powered by a dedicated onboard battery. A Wi-Fi link connects the Jetson to a remote laptop for monitoring and logging, and to a handheld controller for teleoperation. Our software stack (Fig. 2(b)) uses a LiDAR-centric SLAM pipeline [segalGeneralizedICP2009, kümmerleg2o2011] with multi-sensor fusion. Inertial Measurement Unit (IMU) data and legged odometry are fused via an Extended Kalman Filter (EKF) to produce a stable base-frame pose, which is broadcast through the ROS TF tree to align the LiDAR and robot frames. The LiDAR point clouds are processed by the Simultaneous Localization and Mapping (SLAM) front-end with motion compensation and incremental registration guided by the fused odometry. The back-end then optimizes the trajectory and builds a global map, enabling accurate pose estimation.
@@ 3.1.2 Semantic Target Data Collection
The data collected in this work is to support the evaluation of video world models on mapless semantic navigation tasks. The objective is to assess whether a model can infer a plausible motion trend toward a specific semantic target in the scene. Therefore, the robot dog is always initialized with the target already present in the field of view. To guarantee a clear and unambiguous motion trend, a human operator manipulates the robot using a direct goal-approach strategy, maintaining constant forward speed and minimizing unnecessary heading oscillations. This avoids exploration and results in stable and reproducible motion patterns across different scenes.
Each processed data sample in the Target dataset consists of four components: a video sequence, a ground truth trajectory, a semantic target specified in the form of a prompt, and a point cloud map. Video frames are captured at 25Hz, yielding approximately 10 seconds of continuous observation per sample. The first 2 seconds are used for scale factor computation in Sec. 3.2.1, the latter 8 seconds are used as future frames for model inference and training. Each prompt is a concise imperative command expressed in natural language, ensuring that the navigation outcome can be judged with respect to a single well-defined goal in the scene. The semantic targets are pre-selected by human experts to ensure diversity and relevance for navigation tasks. Each target is defined as either explicit or implicit. Explicit targets specify the object name directly, whereas implicit targets describe the object through its attributes or functions embedded in the prompt context (Fig. 4). Among the 450 collected samples, 232 are indoors and 218 are outdoors, 72 have implicit targets and 378 have explicit targets. 125 samples are curated as the benchmark set based on scene/target diversity and a balanced distribution of left/forward/right motion trends, while the remaining 325 samples are used exclusively as training set. Fig. 3 demonstrates the benchmark dataset structure.
@@ 3.2 Target Benchmark Architecture
The Target benchmark (Fig. 5) consists of two main components designed to quantify world model planning capabilities. First, the world-decoder extracts camera trajectories from generated videos using state-of-the-art 3D reconstruction methods. Second, the path evaluation module computes a score focusing on approach tendency using a comprehensive suite of metrics.
Table 2: Comparison of Video World Model Benchmarks. Target-Bench provides invariant, task-driven evaluation of world-model planning in real-world robot scenes, and includes tasks that require implicit semantic reasoning, without relying on subjective model-based scoring, thus reducing judge uncertainty.
Benchmark Task- Driven Real-Rob. Scene Invari. Eval. Visual Consist. Spatial Estim. Implicit Reasoning
VBench [huang2024vbench]
[…]
EWMBench [yue2025ewmbench]
WorldEval [li2025worldeval]
[…]
World-in-World [zhang2025world]
[…]
RBench [deng2026rethinking]
Target-Bench
@@ 3.2.1 World-Decoder
As shown in Fig. 5, to extract camera trajectories from WM’s generated videos, we employ three state-of-the-art 3D reconstruction methods: VGGT [wang2025vggtvisualgeometrygrounded], SpaTracker [xiao2025spatialtrackerv2], and ViPE [huang2025vipe]. Each method processes video frames to reconstruct camera poses.
VGGT. Visual Geometry Grounded Transformer (VGGT) is a feed-forward transformer that directly predicts camera poses, depth maps, and point clouds from multi-view images. Given a sequence of $S$ images $\mathcal{I}=\{I_{1},I_{2},\ldots,I_{S}\}$, VGGT first encodes them through a vision transformer to obtain aggregated tokens. A camera head then predicts a pose encoding $\mathbf{p}_{s}=[\mathbf{T}_{s};\mathbf{q}_{s};\mathrm{\mathbf{fov}}_{s}]\in\mathbb{R}^{9}$ for each frame $I_{s}$, where $\mathbf{T}_{s}\in\mathbb{R}^{3}$ is the camera translation vector, $\mathbf{q}_{s}\in\mathbb{R}^{4}$ is the rotation quaternion, and $\mathrm{\mathbf{fov}}_{s}\in\mathbb{R}^{2}$ represents the horizontal and vertical field of view. This encoding is converted to standard camera parameters $\mathbf{E}_{s}=[\mathbf{R}(\mathbf{q}_{s})\mid\mathbf{T}_{s}]\in\mathbb{R}^{3\times 4}$, where $\mathbf{R}(\mathbf{q}_{s})$ converts the quaternion to a rotation matrix.
SpaTracker. SpaTracker extends VGGT with tracking for better temporal consistency. It adopts a two-stage design: (1) VGGT4Track predicts initial poses $\{\mathbf{E}_{s}^{(0)}\}_{s=1}^{S}$ and depths $\{D_{s}\}_{s=1}^{S}$; (2) a tracking module refines these poses via point correspondences across frames. Given query points $\mathbf{Q}\in\mathbb{R}^{N\times 3}$ on the first frame, the tracker optimizes camera poses through bundle adjustment: $\{\mathbf{E}_{s}^{*}\}=\arg\min_{\{\mathbf{E}_{s}\}}\sum_{s=1}^{S}\sum_{n=1}^{N}\rho\left(\|\pi(\mathbf{K}_{s}\mathbf{E}_{s}\mathbf{X}_{n})-\mathbf{x}_{n,s}\|^{2}\right)$, where $\mathbf{X}_{n}$ is the 3D position of point $n$, $\pi(\cdot)$ the projection, $\mathbf{x}_{n,s}$ the tracked 2D point, and $\rho(\cdot)$ a robust loss. The result is a refined camera-to-world trajectory $\{\mathbf{C}_{2W,s}\}_{s=1}^{S}$, where each $\mathbf{C}_{2W,s}\in\mathbb{R}^{4\times 4}$ is the full camera-to-world transform.
ViPE. ViPE is a SLAM-based visual-inertial pipeline providing metric-scale poses by fusing visual and inertial data. Given RGB frames and IMU inputs, it performs incremental pose estimation with loop closure. Unlike VGGT and SpaTracker, ViPE directly outputs paths in metric units via sensor fusion without requiring scale recovery. The final path is a sequence of SE(3) transformations $\{\mathbf{T}_{s}\}_{s=1}^{S}$, where each $\mathbf{T}_{s}\in\mathbb{R}^{4\times 4}$ encodes rotation and translation in meters.
@@
Monocular methods such as VGGT and SpaTracker estimate camera motion only up to an unknown scale. We restore metric consistency at the segment level by anchoring predictions to a single scalar scale factor $\lambda$ derived from real displacement $d_{\mathrm{real}}$ using the first and last ones from the scale factor recovery frames that result in two predicted extrinsic matrices $\mathbf{E}_{f},\mathbf{E}_{l}\in\mathbb{R}^{3\times 4}$ which are corresponding to the two predicted camera poses enclosed by green dotted box in Fig. 5. We extract the translation vectors $\mathbf{t}_{f}$ and $\mathbf{t}_{l}$ as the fourth column of $\mathbf{E}_{f}$ and $\mathbf{E}_{l}$. The predicted displacement is $d_{\mathrm{pred}}=\|\mathbf{t}_{f}-\mathbf{t}_{l}\|_{2}$, and the scale factor is $\lambda=d_{\mathrm{real}}/d_{\mathrm{pred}}$. Then for all other predicted future frames, we rescale their predicted translation vectors together, $\mathbf{t}_{s}^{\mathrm{scaled}}=\lambda\,\mathbf{t}_{s}$ for $s=1,\ldots,S$ with $S$ be the total number of all frames. This lightweight mechanism preserves relative geometry while lifting trajectories to metric values, enabling quantitative comparison.